# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 3.2 Hybrid variational Bayes
To efficiently handle missing values $\textbf{y}_{u}$ in SEMs described in Section 2, we extend the hybrid VB method of Wijayawardhana et al., (2025) that uses the Gaussian variational approximation with a factor covariance structure to approximate the posterior distribution of the parameters, while employing an MCMC sampler to generate samples of $\textbf{y}_{u}$ from its conditional distribution, given $\textbf{y}_{o}$, m, $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. We first factorise the variational distribution that approximates $\boldsymbol{\xi},\boldsymbol{\psi}$, and $\textbf{y}_{u}$, $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ as follows:
@@
where $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ represents the conditional distribution of $\textbf{y}_{u}$, given the observed data ($\textbf{y}_{o}$, m), and the model parameters ($\boldsymbol{\xi}$ and $\boldsymbol{\psi}$). The term $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ is the Gaussian variational approximation of Ong et al., (2018), used to approximate the posterior distribution of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. We rewrite the ELBO in Equation (S6) by substituting the factorised variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ from Equation (15). Using Bayes’ rule, it is straightforward to show
@@
$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}p({\textbf{y}}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda}),$ (16)
@@
where $\mathcal{L}^{0}(\boldsymbol{\lambda})$ is the ELBO resulting from approximating only the posterior distribution of parameters $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, $p(\boldsymbol{\xi},\boldsymbol{\psi}\mid\textbf{y}_{o},\textbf{m})$, directly using the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$; see Wijayawardhana et al., (2025) for the complete proof.
@@
We assume that the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, follows a multivariate normal distribution with a factor covariance structure (Ong et al.,, 2018), $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu}_{s_{m}},\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2}\right)$, where $\boldsymbol{\mu}_{s_{m}}$ is an $s_{m}\times 1$ vector of variational means, $\textbf{B}_{s_{m}}$ is an $s_{m}\times p$ matrix with the upper triangular elements set to zero, and $\textbf{D}_{s_{m}}$ is an $s_{m}\times s_{m}$ diagonal matrix with positive diagonal elements $\textbf{d}_{s_{m}}=(d_{1},\dots,d_{s_{m}})^{\top}$. For all SEMs with missing values, the value of $s_{m}$ is provided in Table 2. The vector of variational parameters is given by $\boldsymbol{\lambda}_{s_{m}}=\left(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top}\right)^{\top}$, where the ’vech’ operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the superdiagonal elements) of the matrix.
@@
In the SGA algorithm (Equation (14)), we compute an unbiased estimate of the ELBO gradient, $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\lambda)$, denoted as $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\lambda)}$, at each iteration using the reparameterisation trick (Kingma and Welling,, 2013). To do this, we first generate samples from the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ by drawing ${\boldsymbol{\delta}^{0}}=(\boldsymbol{\eta}^{0},\boldsymbol{\epsilon}^{0})\sim N(\mathbf{0},\mathbf{I}_{s_{m}+p})$, where $\boldsymbol{\eta}^{0}$ is a $p$-dimensional vector and $\boldsymbol{\epsilon}^{0}$ is an $s_{m}$-dimensional vector. Next, we compute $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$. The density of $\boldsymbol{\delta}^{0}$ is denoted by $f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})$. Let $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, where the product density is expressed as $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}}))$. Finally, the transformation from $\boldsymbol{\delta}$ to the parameter and augmented missing value space is given by $((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top},{\textbf{y}}_{u}^{\top})^{\top}=t(\boldsymbol{\delta},\boldsymbol{\lambda}_{s_{m}})=(t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top},\textbf{y}_{u}^{\top})^{\top}=((\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0})^{\top},\textbf{y}_{u}^{\top})^{\top}$.
@@
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}))\right],$ (17)
where $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\lambda_{s_{m}}}}$ is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}_{s_{m}}=(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top})^{\top}$; see Section S4.1 of the online supplement for the proof. The expressions for $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}$ and $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ can be found in Section S4.1 of the online supplement. In addition, the expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ are provided in Section S4.3 for different SEMs.
Algorithm 1 outlines the HVB algorithm. For all four SEMs with missing data, the conditional distribution of missing data; $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ is not available in closed form. To sample from $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ in step 5 of Algorithm 1, we use the Markov chain Monte Carlo (MCMC) steps outlined in Algorithm 2.
@@
1:  Initialise $\boldsymbol{\lambda}_{s_{m}}^{(0)}=(\boldsymbol{\mu}_{s_{m}}^{\top(0)},\textrm{vech}{(\textbf{B}_{s_{m}})}^{\top(0)},\textbf{d}_{s_{m}}^{\top(0)})^{\top}$ and set $t=0$
@@
3:     Generate $({\boldsymbol{\eta}^{0}}^{(t)},{\boldsymbol{\epsilon}^{0}}^{(t)})\sim N(\textbf{0},\textbf{I}_{s_{m}+p})$
4:     Generate $(\boldsymbol{\xi}^{(t)^{\top}},\boldsymbol{\psi}^{(t)^{\top}})^{\top}\sim q_{\boldsymbol{\lambda}_{s_{m}}^{(t)}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ using its reparameterised representation.
@@
6:     Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}{(\textbf{B}_{s_{m}})}}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}\widehat{{\nabla_{\textbf{d}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S44), (S45) and  (S46) in Section S4.1 of the online supplement.
[…]
11:     Set $\boldsymbol{\lambda}_{s_{m}}^{(t+1)}=(\boldsymbol{\mu}_{s_{m}}^{\top(t+1)},\textrm{vech}(\textbf{B}_{s_{m}})^{\top(t+1)},\textbf{d}_{s_{m}}^{\top(t+1)})^{\top}$, and $t=t+1$
@@
5:     Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{{i-1}},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\widetilde{\textbf{y}}_{u}}^{\top})$ and $\textbf{y}_{i-1}=(\textbf{y}^{\top}_{o},{\textbf{y}^{\top}_{u,{i-1}}})$
@@
3.3 Bayesian model comparison
To assess model fit in both the simulation study (Section 4) and the real-world analysis (Section 5), we employ extended versions of the standard Deviance Information Criterion (DIC, Spiegelhalter et al., (2002)) proposed by Celeux et al., (2006). This section first outlines DIC methods for models without missing data, then for those with missing data.

### v2

@@ 4.2 Hybrid variational Bayes
This section extends the hybrid VB (HVB) algorithm of Wijayawardhana et al., (2025) to efficiently handle missing values $\textbf{y}_{u}$ in non-Gaussian SEMs proposed in Section 2. The approach employs a Gaussian variational approximation with a factor covariance structure (Ong et al.,, 2018) to approximate the posterior distribution of the parameters, while employing an MCMC sampler to generate samples of $\textbf{y}_{u}$ from its conditional distribution, given $\textbf{y}_{o}$, m, $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. We first factorise the joint variational distribution that approximates $\boldsymbol{\xi},\boldsymbol{\psi}$, and $\textbf{y}_{u}$, $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ as follows:
@@
where $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ represents the conditional distribution of $\textbf{y}_{u}$, given the observed data ($\textbf{y}_{o}$, m), and the model parameters ($\boldsymbol{\xi}$, $\boldsymbol{\psi}$). The term $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ is the variational approximation used to approximate the posterior distribution of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. The ELBO in Equation (S6) can be rewritten by substituting the factorised variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ from Equation (18). By applying Bayes’ rule and simplifying, we obtain
@@
$\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{{q}_{\boldsymbol{\lambda}}}\left[\text{log}~p({\textbf{y}}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda}),$ (19)
@@
where $\mathcal{L}^{0}(\boldsymbol{\lambda})$ is the ELBO resulting from approximating only the posterior distribution of parameters ($\boldsymbol{\xi}$, $\boldsymbol{\psi}$), that is, $p(\boldsymbol{\xi},\boldsymbol{\psi}\mid\textbf{y}_{o},\textbf{m})$, directly using the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$; see Wijayawardhana et al., (2025) for the complete proof.
@@
We assume that the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, follows a multivariate Gaussian distribution with a factor covariance structure (Ong et al.,, 2018), $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2}\right)$}, where $\boldsymbol{\mu}$ is an $s\times 1$ vector of variational means, B is an $s\times p$ matrix with the upper triangular elements set to zero ($p<<s$), and D is an $s\times s$ diagonal matrix with positive diagonal elements $\textbf{d}=(d_{1},\dots,d_{s})^{\top}$. For all SEMs with missing values, the value of $s$ is provided in Table 2. The vector of variational parameters is given by $\boldsymbol{\lambda}=\left(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top}\right)^{\top}$, where the ’vech’ operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the super-diagonal elements) of the matrix.
@@
We now describe how the SGA method is used in the proposed HVB algorithm. At each iteration, an unbiased estimate of the ELBO gradient, $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$, denoted $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, needs to be computed for the SGA update given in Equation (17). This is achieved using the reparameterisation trick (Kingma and Welling,, 2013). First, samples are generated from the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ by drawing ${\boldsymbol{\delta}^{0}}=(\boldsymbol{\eta}^{0},\boldsymbol{\epsilon}^{0})\sim N(\mathbf{0},\mathbf{I}_{s+p})$, where $\boldsymbol{\eta}^{0}$ is a $p$-dimensional vector and $\boldsymbol{\epsilon}^{0}$ is an $s$-dimensional vector. Next, we compute $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$. The density of $\boldsymbol{\delta}^{0}$ is denoted by $f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})$. Let $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, where the product density is expressed as $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}))$}. Finally, the transformation from $\boldsymbol{\delta}$ to the parameter and augmented missing value space is given by $((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top},{\textbf{y}}_{u}^{\top})^{\top}=t(\boldsymbol{\delta},\boldsymbol{\lambda})=(t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top},\textbf{y}_{u}^{\top})^{\top}=((\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0})^{\top},\textbf{y}_{u}^{\top})^{\top}$.
@@
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}))\right],$ (20)
where $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$; see Section S5.1 of the online supplement for the proof. The expressions for $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}$ and $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ can be found in Section S5.2 of the online supplement. In addition, the expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ are provided in Section S5.3 of the online supplement for different SEMs.
Algorithm 1 outlines the HVB algorithm. For all four SEMs with missing data, the conditional distribution of missing data, $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ is not available in closed form. To sample from $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ in step 5 of Algorithm 1, we use the MCMC steps outlined in Algorithm 2.
@@
3:  Generate $({\boldsymbol{\eta}^{0}}^{(t)},{\boldsymbol{\epsilon}^{0}}^{(t)})\sim N(\textbf{0},\textbf{I}_{s+p})$
4:  Generate $(\boldsymbol{\xi}^{(t)^{\top}},\boldsymbol{\psi}^{(t)^{\top}})^{\top}\sim q_{\boldsymbol{\lambda}^{(t)}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ using its reparameterised representation.
@@
6:  Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}{(\textbf{B})}}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}\widehat{{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S46), (S47) and  (S48) in Section S5.2 of the online supplement.
7:  Set adaptive learning rates $\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}$ , $\mathbcal{a}^{(t)}_{\text{vech}(\textbf{B})}$ and $\mathbcal{a}_{\textbf{d}}^{(t)}$, using ADADELTA described in Section S3 of the online supplement.
8:  Set $\boldsymbol{\mu}^{(t+1)}=\boldsymbol{\mu}^{(t)}+\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}\circ\widehat{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
9:  Set $\textrm{vech}(\textbf{B})^{(t+1)}=\textrm{vech}(\textbf{B})^{(t)}+\mathbcal{a}_{\text{vech}(\textbf{B})}^{(t)}\circ\widehat{\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
10:  Set $\textbf{d}^{(t+1)}=\textbf{d}^{(t)}+\mathbcal{a}_{\textbf{d}}^{(t)}\circ\widehat{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
11:  Set $\boldsymbol{\lambda}^{(t+1)}=(\boldsymbol{\mu}^{\top(t+1)},\textrm{vech}(\textbf{B})^{\top(t+1)},\textbf{d}^{\top(t+1)})^{\top}$, and $t=t+1$
@@
5:  Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{{i-1}},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\widetilde{\textbf{y}}_{u}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}^{\top}_{o},{\textbf{y}^{\top}_{u,{i-1}}})^{\top}$
@@
5 Bayesian model comparison
To assess the model fit of different SEMs in the subsequent sections of this study, we employ extended versions of the standard Deviance Information Criterion (DIC; Spiegelhalter et al., (2002)) as proposed by Celeux et al., (2006). This section first outlines DIC methods for models without missing data, followed by those for models with missing data.
5.1 Model comparison with full data